2x+4(4x-3)=-3-2(-2x+1)
How do I solve this question?

Answer by gsmani_iyer(201)


2x + 16x -12 = -3 +4x -2 ........ expanded by removing the brackets

= 18x - 12 = 4x - 5

= 18x - 4x = 12 - 5 ............by bringing variables oneside and numbers other side
= 14x = 7 ..........by simplification

= x = 7/14

= x = 1/2 or 0.5 Answer

gsmani9


=